### v3

# An Efficient Learning Control Framework with Sim-to-Real for String-Type Artificial Muscle-Driven Robotic Systems

Abstract
Robotic systems driven by artificial muscles present unique challenges due to the nonlinear dynamics of actuators and the complex designs of mechanical structures. Traditional model-based controllers often struggle to achieve desired control performance in such systems. Deep reinforcement learning (DRL), a trending machine learning technique widely adopted in robot control, offers a promising alternative. However, integrating DRL into these robotic systems faces significant challenges, including the requirement for large amounts of training data and the inevitable sim-to-real gap when deployed to real-world robots. This paper proposes an efficient reinforcement learning control framework with sim-to-real transfer to address these challenges. Bootstrap and augmentation enhancements are designed to improve the data efficiency of baseline DRL algorithms, while a sim-to-real transfer technique, namely randomization of muscle dynamics, is adopted to bridge the gap between simulation and real-world deployment. Extensive experiments and ablation studies are conducted utilizing two string-type artificial muscle-driven robotic systems including a two degree-of-freedom robotic eye and a parallel robotic wrist, the results of which demonstrate the effectiveness of the proposed learning control strategy.
Index Terms:
Reinforcement learning for robot control, artificial muscle, sim-to-real, sample efficiency.
I Introduction
Artificial muscles, designed to mimic biological counterparts, generate mechanical forces and movements in response to external stimuli like electrical or chemical signals. Well known for their high power-to-weight ratios and inherent compliance[1], artificial muscles offer a unique combination of strength and flexibility conducive to versatile and adaptive robot designs. Over the past decade, robots driven by artificial muscles have achieved significant advancements, demonstrating impressive performance across diverse domains, including but not limited to biomimetic robots[2, 3], robot manipulators[4, 5], soft robots[6, 7], medical robot[8], origami robots[9], and marine robots[10].
Previous research efforts have primarily centered around structural design and open-loop motion performance evaluation[2, 3, 4, 5, 6, 7, 8, 9, 10]. Despite their potential, achieving accurate control of string-type artificial muscle-driven robots remains challenging due to their highly nonlinear and complex dynamics. Traditional control strategies, such as proportional-integral-derivative (PID) control[11, 12, 13], energy-based nonlinear control[14], and sliding mode control[15, 16] have been investigated to mitigate these challenges. Although these approaches have demonstrated success in specific tasks, a comprehensive, unified control strategy for string-type artificial muscle-driven robots is still lacking. As deep reinforcement learning (DRL) rapidly develops, learning-based controllers[17, 18, 19, 20, 21] have emerged as a promising alternative, offering the potential to overcome the limitations of conventional methods. The standard procedure of applying DRL in the targeted robotic systems typically follows the schematic illustrated in Fig. 1. Initially, identifying the relevant parameters and establishing a simulation environment enable a safe and efficient training process. The DRL algorithms are then trained within this simulation environment to develop an effective control policy. Finally, the learned control policy is transferred to the physical robotic platform to achieve the desired control objectives.
Figure 1: Schematic of standard procedure of applying DRL control in artificial muscle-driven robotic systems.
However, several challenges persist in the applications of DRL to string-type artificial muscle-driven robots. One major issue is the sim-to-real gap, which arises from inaccuracies in dynamic models, system identification errors, and variability in muscle string parameters[22, 23]. These factors often result in control strategies that demonstrate strong performance in simulation but struggle to transfer effectively to real-world applications. Another challenge is the low learning efficiency of DRL algorithms, which require substantial amounts of training data and iterations due to the complexity of neural networks and the trial and error learning process[24]. Addressing these two challenges is crucial for advancing control technologies in artificial muscle-driven robots, which are increasingly employed in soft robotics and bio-inspired systems.
This paper presents a DRL-based control strategy designed to enhance the data efficiency and real-world applicability of controllers for artificial muscle-driven robots. The proposed framework is evaluated on two robotic systems actuated by artificial muscle strings: a biomimetic robotic eye driven by super-coiled polymers (SCPs)[25, 26] and a parallel robotic wrist actuated by twisted-coiled actuators (TCAs)[27, 28]. Specifically, we introduce a novel augmentation method for state-based data, which transforms a single transition data tuple into several, and we leverage an easy-to-implement PID controller to generate trajectories for early-stage training of DRL agents. Compared to the baseline algorithm (i.e., Soft Actor Critic [29]), the required training data is reduced by approximately $60\%$ in the targeted robotic systems. Additionally, we leverage a zero-shot sim-to-real transfer technique on artificial muscle dynamics. The experimental results demonstrate a significant improvement in real-world control performance with the averaged error reduced by $36.3\%$ through the application of this method.
The remainder of this paper is organized as follows. Section II provides a review of the relevant literature. Section III introduces the basic concepts of applying reinforcement learning to the robotic systems studied. Section IV offers a concise overview of the robotic platforms utilized in this study. In Section V, we present a detailed description of our proposed efficient learning control framework. Section VI and Section VII discuss the experimental results, including training outcomes in simulation and control performances in real-world scenarios, respectively. Finally, concluding remarks and current limitations are discussed in Section VIII.
II Related Work
The control of artificial muscle-driven robotic systems using learning-based methods has garnered significant attention in recent years[30, 31, 20, 19, 32, 21, 18, 33]. Among these approaches, DRL has demonstrated effectiveness in handling diverse and complex tasks, positioning it as a promising control solution for a wide range of applications. An increasing number of researchers have leveraged DRL algorithms to address control challenges in artificial muscle-driven robots, spanning applications ranging from dexterous grasping of robotic hands[17] to agile motion control of soft robotic manipulators[33, 18], and other related applications[32, 21]. Many of these studies have noted that the low sample efficiency often leads to prolonged training time[20, 21], but few have focused specifically on improving training efficiency. Several existing state-of-the-art (SOTA) reinforcement learning (RL) methods may be applicable[34, 35, 36, 37], but they have not been evaluated within the context of the robotic systems we are addressing. The unique nonlinear and slow response characteristics of artificial muscle may present challenges that are not adequately addressed by these existing methods. Our previous work[19] explored low learning efficiency problem but was limited to simulation-based experiments.
Due to the sim-to-real gap, control strategies that perform well in simulation may not seamlessly transfer to real-world applications[33, 18]. Despite being labor-intensive, some researchers[31, 32] have attempted to directly train DRL algorithms on real robots. However, the artificial muscles considered in this paper are all thermally actuated, which inherently leads to energy inefficiency and limited cycling rates[38]. Continuous rapid transitions between heating and cooling often cause irreversible damage to the muscles, making direct training on real robots unsafe. It is worth noting that in recent years, there have been some efforts to bridge the sim-to-real gap of artificial muscles. Many researchers[22, 23, 39] have proposed more accurate yet significantly more complex dynamic models. In this paper, however, we aim to directly improve the performance of DRL-based controllers by employing a sim-to-real transfer technique, thereby reducing the need for complicated mathematical modeling and precise parameter identification for artificial muscles.
III Preliminaries
In this section, we model the control problem of targeted robotic systems as a reinforcement learning problem by Markov Decision Process (MDP), which consists of the tuple $(\mathcal{S},\mathcal{A},\mathcal{R},f,\gamma)$.
  • $\mathcal{S}$ is the state space. The state vector $s_{t}\in\mathcal{S}$ generally includes the current state variables of the robotic system and the target positions and/or orientations. In this paper, the motion state variables of the robotic systems are collectively defined as vector $\bm{x}$. The system output vector, containing the variables to be controlled, is denoted by $\bm{y}$. Correspondingly, the target system output is denoted by $\bm{y}^{*}$. In this way, the state vector $s_{t}$ is represented as $s_{t}=(\bm{x}_{t},\bm{y}^{*})$.
  • $\mathcal{A}$ is the action space, where each action $a_{t}\in\mathcal{A}$ is generated by the DRL agent based on the current state $s_{t}$ and its own policy.
  • $\mathcal{R}$ is the reward function that maps the current state $s_{t}$ and action $a_{t}$ to an immediate reward $r_{t}=\mathcal{R}(s_{t},a_{t})$. The design of reward function plays a vital role in the learning process of the DRL agents, which learn the optimal policy under the guidance thereby. In this paper, we develop a generalized quadratic-form reward function designed for the robotic systems of interest:
    $r_{t}=-(e^{\top}_{t}Q_{e}e_{t}+a^{\top}_{t}R_{a}a_{t})+r_{\text{bonus}},$ (1)
    where $e_{t}=|\bm{y}^{*}-\bm{y}_{t}|$ is the system tracking error, and $Q_{e}$ and $R_{a}$ represent the output and control weight matrices, respectively. A bonus reward $r_{\text{bonus}}$ is given when the system tracking error falls below a sufficiently small threshold.
  • $f$ represents the state transition function and $\gamma$ is the discount factor, which is set to $0.99$ in our experiments.
Figure 2: Illustrations of SCP and TCA muscles. (a) Crimp fixtures for the SCP muscle ends (left), and a close-up shot of the fabricated SCP muscle (right); and (b) conductive ring with SMA wire wrapped around for the TCA muscle ends (left), and a close-up shot of the fabricated TCA muscle (right).
IV Experimental Platforms
This section presents an overview of the actuators and robotic platforms used in this paper, including SCP and TCA artificial muscles (Fig. 2), a 2-DOF robotic eye (Fig. 3) and a parallel robotic wrist (Fig. 4). For each system, we briefly describe the hardware design, state space, action space, and reward function.
IV-A Artificial Muscles
Numerous fabrication methods for SCP and TCA muscles have been investigated in the literature[25, 26, 27, 28]. Specifically, for our robotic eye design, the SCP actuator is fabricated by sequentially twisting, coiling, and double-backing using a 145 cm long silver-plated conductive nylon thread (Statex’s Shieldex 117/17 dtex 2 ply HC+B). As illustrated in Fig. 2(a), the ends of the SCP muscles are crimped using fold-over silver/copper bead crimps. They are connected to N-Channel enhancement MOSFETs (FQP30N06L) which provide electric power to the muscles. Application of voltage induces Joule heating on the actuator which causes the coils to expand laterally, thus generating contracting force.
Regarding our TCA muscles, the Shape Memory Alloy (SMA) wires (0.18 mm diameter) serve as the conductive threads and skeletons. Highly elastic spandex fibers (creora spandex, power fit, 840 denier) are tightly wound around the SMAs, and the resulting composite fibers are then twisted into a coil configuration. As shown in Fig. 2(b), both ends of the SMA wire are wrapped around conductive rings, which are also connected to MOSFETs. When voltage potential is applied to the rings, the skeleton generates Joule heat and the surrounding spandex fibers will contract in longitudinal direction, which leads to the untwisting of the compound twisted fiber structure. The TCA’s coiled configuration converts this untwisting motion into linear contraction. Additionally, the SMA wire not only provides Joule heating, but also contributes additional contraction force owing to its shape memory effect.
Figure 3: The 2-DOF robotic eye prototype driven by two pairs of SCP muscles.
IV-B Robotic Eye
The high occurrence of unsuccessful corrective surgeries for ocular motor disorders underscores the need for a muscle-driven robotic eye platform to assist ophthalmologists in gaining deeper insights into the biomechanics of human eye movement. In our previous work[25], we proposed a prototype of a 2-DOF robotic eye and derived the geometric and dynamic models specific to this prototype. As illustrated in Fig. 3, the hemispherical eyeball is driven by two pairs of SCP muscle actuators, accommodating a Raspberry Pi Mini Camera (RPiMC). The primary control objective is to precisely and rapidly adjust the eyeball’s orientation towards the desired target angles by regulating the SCP muscles, namely foveation control in the eye movement study.
The eyeball’s center of rotation, denoted as $\bm{O}\in\mathbb{R}^{3}$, remains fixed at the origin within the reference frame $O_{\mathrm{xyz}}$. Here, the axes $\mathrm{x}$, $\mathrm{y}$, and $\mathrm{z}$ correspond to the transverse, sagittal, and vertical axes of the eyeball, respectively. Rotations around these axes are expressed as pitch $\theta$, roll $\phi$, and yaw $\psi$. The robotic eye system’s state variables are represented as
$\bm{x}=[\theta,\dot{\theta},\psi,\dot{\psi},\Delta T_{1},\Delta T_{2},\Delta T_{3},\Delta T_{4}]^{\top},$ (2)
where $\theta,\dot{\theta},\psi,\dot{\psi},$ represent the pitch, yaw angles, and their respective angular velocities for control. $\Delta T_{i}$ represents the temperature difference between the ambient temperature $T_{0}$ and the current temperature $T_{i}$ of muscle $m_{i}$.
Figure 4: The parallel robotic wrist prototype actuated by three TCA muscles.
Focusing primarily on the robotic eye’s orientation, the system’s output vector is denoted as $\bm{y}=[\psi,\dot{\psi},\theta,\dot{\theta}]^{\top}$. Due to the difficulty and poor accuracy of measuring the temperature of each muscle in real time, the state vector of the DRL agent is defined by
$s_{t}=[\theta_{t},\dot{\theta}_{t},\psi_{t},\dot{\psi}_{t},\theta^{*},\psi^{*}]^{\top}.$ (3)
Here, $\theta^{*}$ and $\psi^{*}$ represent the target angles. The Euler angles and corresponding angular velocities are measured through visual feedback from the embedded RPiMC. The actuating voltage applied to muscle $m_{i}$ is denoted as $V_{i}$. Given the antagonistic configuration of the muscle pairs, the control input is defined by a constrained action vector, i.e.,
$a_{t}=[$ $-V_{1}+V_{2},-V_{3}+V_{4}]^{\top}$ (4)
[…]
The constraint requires that only one muscle is activated at each time step in the pairs $(m_{1},m_{2})$ and $(m_{3},m_{4})$. The reward function follows the structure of (1), where the output and control weight matrices were set to $Q_{e}=\text{diag}\{0.05,0.25,0.05,0.25\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively. They were chosen based on empirical tuning to balance the trade-off between output error and control effort. The bonus reward is granted when the tracking error is sufficiently small, given by
$r_{\text{bonus}}=2\times\left(\mathds{1}\{|\psi-\psi^{*}|<0.3\}+\mathds{1}\{|\theta-\theta^{*}|<0.3\}\right).$ (5)
Here, $\mathds{1}\{\cdot\}$ denotes the binary indicator function that yields a value of $1$ if the given condition holds true, and 00 otherwise.
IV-C Robotic Wrist
Research into artificial muscle-driven robotic wrist joints is of great importance to several domains, including the development of dexterous robotic systems with human-like capabilities, advancements in rehabilitation and prosthetics technology, and the enhancement of robotic manipulation versatility across various fields. We designed a parallel robotic wrist prototype driven by three equally spaced compound TCAs[27] as illustrated in Fig. 4. The base plate remains stationary, while the upper plate, also known as the end effector, moves via three muscle strings. Linkages between the two plates are constructed to support a two degree-of-freedom motion of the upper plate[40]. Given this configuration, we developed a simplified dynamics model under the assumption of a fixed rotational center for the upper plate. Similar to the foveation control of the robotic eye, the control objective entails precise regulation of the moving upper plate’s orientation. This necessitates adjusting the applied actuating voltages across the three muscle strings to drive the upper plate from its initial equilibrium state to predefined target angles.
The reference frame $O_{\mathrm{xyz}}$ is fixed on the moving upper plate, with the $\mathrm{z}$-axis perpendicular to the upper plate and pointing outward. $\theta,\phi,$ and $\psi$ represent the pitch, roll, and yaw angles, respectively. Similar to the robotic eye, the system state variables are defined as
$\bm{x}=[\theta,\dot{\theta},\phi,\dot{\phi},\Delta T_{1},\Delta T_{2},\Delta T_{3}]^{\top},$ (6)
where $\dot{\theta}$ and $\dot{\phi}$ represent the corresponding angular velocities, and $\Delta T_{i}$ is the temperature difference of muscle $m_{i}$ with respect to the ambient environment, with $i\in\{1,2,3\}$. The system output vector is represented as $\bm{y}=[\theta,\dot{\theta},\phi,\dot{\phi}]^{\top}$. Likewise, the muscle’s temperature is not used as an input for the DRL agent, and thus, the state vector is defined as
$s_{t}=[\theta_{t},\dot{\theta}_{t},\phi_{t},\dot{\phi}_{t},\theta^{*},\phi^{*}]^{\top},$ (7)
where $\theta^{*}$ and $\phi^{*}$ denote the target angles, with the respective Euler angles and angular velocities measured using the inertial measurement unit (IMU) positioned at the center of the upper plate. Additionally, the input action vector $a_{t}$ of the system is defined as:
[…]
The output and control weight matrices of the reward function are set to $Q_{e}=\text{diag}\{0.05,0.2,0.05,0.2\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively. When the tracking error is sufficiently small, a bonus reward is granted, i.e.,
$r_{\text{bonus}}=2\times\left(\mathds{1}\{|\theta-\theta^{*}|<0.5\}+\mathds{1}\{|\phi-\phi^{*}|<0.5\}\right).$ (9)
V Learning Control Framework
Figure 5: Overview of the proposed efficient learning control framework, including a two-stage training process in simulation and real-world application. Stage 1 utilizes bootstrapping with PID control, aiming at enhancing data efficiency. Stage 2 integrates the baseline SAC learning algorithm with state-based data augmentation and randomization of muscle dynamics for expediting training convergence and bridging sim-to-real gap, respectively.
This section provides a comprehensive overview of the proposed learning control framework, as illustrated in Fig. 5. The soft actor-critic (SAC) algorithm serves as the baseline DRL method. To improve data efficiency, we employ a two-stage training strategy. Initially, a standard PID controller is initialized to generate imperfect trajectories for the DRL agent. Subsequently, the agent interacts directly with the simulation environment to update its policy. All transition data collected from the environment undergoes a state-based augmentation strategy before being stored into the replay buffer. To mitigate the sim-to-real gap, we introduce a randomization method on the artificial muscle dynamics at the beginning of each training episode. Eventually, the trained actors (control policies) can be directly implemented in real-world robotic systems to achieve the desired control performance.
V-A Soft Actor-Critic
Soft Actor-Critic (SAC) [29], a state-of-the-art DRL algorithm, is widely adopted as a strong baseline for continuous control tasks in robotics due to its robustness, sample efficiency, and model-free nature. Unlike traditional DRL algorithms, SAC incorporates an entropy-maximizing approach, with its learning objective defined as:
$J(\pi)=\sum_{t=0}^{T}\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}\left[\mathcal{R}(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot|s_{t}))\right],$ (10)
where $\mathcal{H}(\pi(\cdot|s_{t}))$ represents the entropy of the policy $\pi$, and $\alpha$ serves as a temperature parameter. This entropy-regularized objective improves the robustness of DRL and accelerates its training [24]. To process sequential information critical for addressing partial observability issue and implementing dynamics randomization, we integrate a Gated Recurrent Unit (GRU) [41] as the hidden layer in the actor-critic architecture.
V-B Randomization of Muscle Dynamics
TABLE I: Dynamics parameters of randomization and their respective ranges in the robotic eye
Parameter Value Scailing factor Additive term
[…]
$c$ $(\text{N}/^{\circ}\text{C})$ $0.0055$ uniform$(0.85,1.15)$
$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $0.28$ uniform$(0.8,1.2)$
$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.094$ uniform$(0.85,1.15)$
$R$ $(\Omega)$ $20.0$ uniform$(0.9,1.1)$
$\theta,\psi$ noise $(^{\circ})$ normal$(0,0.1)$
$\dot{\theta},\dot{\psi}$ noise $(^{\circ}/\text{s})$ normal$(0,0.05)$
The dynamic model of the artificial muscles considered in this paper is based on the literature[26], where the muscle’s elasticity behavior is linearized to establish a dynamical system comprising a mechanical model characterized by stiffness and damping, given by
$F=k(x-x_{0})+b\dot{x}+c(T(t)-T_{\text{amb}}),$ (11)
where $x$ and $x_{0}$ denote the loaded length and unloaded length of the actuator, respectively, while $k$, $b$, and $c$ denote the mean stiffness, damping, and temperature effect coefficients of the actuator, respectively. $T(t)$ denotes the current actuator temperature, while $T_{\text{amb}}$ stands for the ambient temperature of the environment. The thermoelectric model of the actuator is expressed as:
$C_{\text{th}}\frac{dT(t)}{dt}=P(t)-\lambda(T(t)-T_{\text{amb}}),$ (12)
where $C_{\text{th}}$ represents the thermal mass of the actuator (measured in $\text{Ws}/^{\circ}\text{C}$), $P(t)$ indicates the heat applied to the actuator, $\lambda$ is the absolute thermal conductivity of the actuator with respect to the ambient environment (in $\text{W}/^{\circ}\text{C}$). The applied power, $P(t)$, is controlled through the voltage applied to the muscle, expressed as $P(t)=V(t)^{2}/R$.
TABLE II: Dynamics parameters of randomization and their respective ranges in the robotic wrist
Parameter Value Scailing factor Additive term
[…]
$c$ $(\text{N}/^{\circ}\text{C})$ $0.0707$ uniform$(0.85,1.15)$
$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $3.06$ uniform$(0.8,1.2)$
$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.1189$ uniform$(0.85,1.15)$
$R$ $(\Omega)$ $10.0$ uniform$(0.9,1.1)$
$\theta,\phi$ noise $(^{\circ})$ normal$(0,0.1)$
$\dot{\theta},\dot{\phi}$ noise $(^{\circ}/\text{s})$ normal$(0,0.05)$
The core idea behind dynamics randomization is to introduce randomness into specific dynamic parameters within the simulator during training. This approach allows the DRL agent to learn a policy adaptable to various dynamics conditions, including those encountered in real-world experiments. Specifically, we randomize several critical muscle parameters within the simulated dynamics environment. A distribution $\rho_{\mu}$ of the muscle parameter set $\mu$ is established to encompass the variability of muscle dynamic parameters for dynamics randomization purposes. This approach is expected to be particularly effective in our setup, as artificial muscles are known to exhibit significant variability due to factors such as material degradation, slight manufacturing differences, and environmental temperature changes. According to the muscle models (11) and (12), the dynamic parameter set $\mu$ for randomization is given by
$\mu:=\left\{k,b,c,C_{\text{th}},\lambda,R\right\}.$ (13)
Table I and II outline the distribution $\rho_{\mu}$ of SCP and TCA muscles, respectively, along with the observation noise added for the robotic eye and wrist systems. $\rho_{\mu}$ is centered on the mean values derived from system identification to align the data distribution in simulation more accurately with that in reality. To strike the right balance between randomization and fidelity, we carefully tune the range and distribution of randomized parameters based on empirical observations of muscle behavior in real-world scenarios. Specifically, the range of each parameter is determined based on the distribution observed across different real muscle strings, as outlined in [26]. This approach allows us to maintain a high degree of fidelity while simultaneously enabling the controller to generalize effectively to real-world uncertainties. Details on the impact of varying randomization ranges on experimental results are discussed in Section VII-C.
During the training process, the physical parameters of the actuators are randomized within predefined scaling ranges (i.e., $\mu\sim\rho_{\mu}$) at the start of each training episode and remain constant throughout. Additionally, Gaussian noise is added to the policy’s observation inputs to simulate the noise typically present in real-world sensor data. This observation noise varies at each time step and across different state variables. Although this study employs a simplified dynamic model for artificial muscles, the proposed method can be extended to more complex muscle dynamics models to further enhance control performance.
V-C Bootstrap via PID
On-line DRL agents primarily rely on random exploration [24]. However, due to the slow response characteristics of artificial muscles, random exploration policies often fail to explore the state space comprehensively and efficiently, leading to low sample efficiency. To address this limitation, we accelerate the early-stage training by incorporating demonstration data into the initial replay buffer. For muscle-driven robotic systems, PID controllers are typically sufficient to approach target states and deliver moderate control performance. As such, they can generate trajectories that explore the defined state space more comprehensively, providing a beneficial starting point for the DRL learning process.
Specifically, we initialize a PID controller that offers a barely acceptable (and typically not excellent) control performance before DRL training begins. Within the first $M$ episodes of training, the PID controller interacts with the environment and generates trajectories, which are then stored in the replay buffer along with the rewards. Unlike traditional behavior cloning methods, we do not employ a supervised loss for demonstration learning, allowing the agent to avoid strictly mimicking the PID control policy. Subsequently, the agent continues interacting with the environment to refine its policy, eventually surpassing the PID controller and achieving superior control performance.
V-D Augmentation of State Vector
Data augmentation refers to a class of techniques aimed at increasing the size and improving the quality of training datasets, thereby facilitating the development of more robust deep learning models[42]. The fundamental concept of data augmentation involves transforming limited data into diverse and functionally equivalent datasets. Given the labor-intensive and computationally expensive nature of collecting training data in robotic systems, we integrate data augmentation into the agent’s learning process. This paper proposes a novel and straightforward state-based data augmentation technique designed for our targeted robotic systems.
Algorithm 1 SAC-BAR Learning Control
1:  Initialize critic networks $Q_{\theta_{1}},Q_{\theta_{2}}$, and actor network $\pi_{\phi}$ with random parameters $\theta_{1},\theta_{2},\phi$
2:  Initialize target network parameters $\bar{\theta}_{1}\leftarrow\theta_{1},\bar{\theta}_{2}\leftarrow\theta_{2}$
3:  Initialize an empty replay buffer $\mathcal{B}$
4:  Initialize a PID controller $\pi_{\text{PID}}$
5:  Initialize dynamics parameters distribution $\rho_{\mu}$
6:  for $\text{step}=1$ to $M$ do
7:     Sample dynamics $\mu\sim\rho_{\mu}$
8:     Generate rollout $\tau_{0}=(s_{0},a_{0},r_{0},...,s_{T})$ with dynamics $\mu$ and $a_{t}\sim\pi_{\text{PID}}(s_{t})$
9:     for $i=1$ to $n$ do
10:        Create new rollouts $\tau_{0}\xrightarrow{AUG}\tau_{i}$
11:     end for
12:     Store $\tau_{0},\tau_{1},...,\tau_{n}$ into $\mathcal{B}$
13:  end for
14:
15:  for $\text{step}=M+1$ to $N$ do
16:     Sample dynamics $\mu\sim\rho_{\mu}$
17:     Generate rollout $\tau_{0}=(s_{0},a_{0},r_{0},...,s_{T})$ with dynamics $\mu$ and $a_{t}\sim\pi_{\phi}(s_{t})$
18:     for $i=1$ to $n$ do
19:        Create new rollouts $\tau_{0}\xrightarrow{AUG}\tau_{i}$
20:     end for
21:     Store $\tau_{0},\tau_{1},...,\tau_{n}$ into $\mathcal{B}$
22:     for $i=1$ to $k$ do
23:        Sample a mini-batch of rollouts from $\mathcal{B}$
24:        Calculate the gradients $\nabla J_{Q}(\theta_{i}),\nabla J_{\pi}(\phi),\nabla J(\alpha)$
25:        Update network weights:
26:        $\theta_{i}\leftarrow\theta_{i}-\lambda_{Q}\nabla J_{Q}(\theta_{i})$
27:        $\phi\leftarrow\phi-\lambda_{\pi}\nabla J_{\pi}(\phi)$
28:        $\bar{\theta}_{i}\leftarrow\tau\theta_{i}+(1-\tau)\theta_{i}$
29:        Adjust temperature parameter:
30:        $\alpha\leftarrow\lambda_{\alpha}\nabla J(\alpha)$
31:     end for
32:  end for
As commonly acknowledged, rewards serve as pivotal guides in the learning process of DRL agents. Therefore, our objective is to generate additional data transitions with different rewards. As described in Section III, the reward $r_{t}$ is calculated as:
$r_{t}=\mathcal{R}(\bm{x}_{t},\bm{y}^{*},a_{t}).$ (14)
Here, $\bm{x}_{t}$ and $a_{t}$ are directly related to system dynamics, whereas the target output vector $\bm{y}^{*}$ is manually set and changeable. Consequently, the proposed augmentation of state vectors proceeds as follows. At each time step $t$, we fix the motion state vectors $\bm{x}_{t},\bm{x}_{t+1}$ and the action $a_{t}$, and vary the target vector $\bm{y}^{*}$. In this paper, we introduce a random additive term into $\bm{y}^{*}$ as follows
$\bm{y}^{*}_{\text{new}}=\bm{y}^{*}\pm\delta\bm{Z}.$ (15)
Here, $\bm{y}^{*}_{\text{new}}$ denotes the new target vector. $\delta$ acts as a scalar factor to control the scope of augmentation, which largely depends on the defined range. $\bm{Z}$ is a random vector sharing identical dimensions with $\bm{y}^{*}$, wherein the elements of $\bm{Z}$ range from 00 to $1$. Thus, we derive a pair of new states $s^{\prime}_{t}$ and $s^{\prime}_{t+1}$. After that, a new corresponding reward $r^{\prime}_{t}$ is calculated using (14). Following the above steps, the original transition tuple is transformed into a new one, i.e.,
$(s_{t},a_{t},r_{t},s_{t+1})\xrightarrow{AUG}(s^{\prime}_{t},a_{t},r^{\prime}_{t},s^{\prime}_{t+1}),$ (16)
which creates different transition tuples at a negligible computational cost.
In this paper, we denote the bootstrap method by the letter ’B’, augmentation method by the letter ’A’, and the randomization of muscle dynamics by the letter ’R’. Thus, our proposed SAC algorithm with all three improvements is denoted by SAC-BAR. The pseudo code of our overall SAC-BAR learning control framework is presented in Algorithm 1.
Figure 6: Reward trajectories with respect to the training episode number. (a) Comparison results between our SAC-BAR and SOTA methods including REDQ, DroQ, RLPD, and X-QL. (b) Results of ablation study where ‘B’, ‘A’, and ‘R’ indicate bootstrap, augmentation, and randomization techniques, respectively.
VI Simulation Experiments
In this section, we present the training results from simulation environments, comparing our proposed SAC-BAR against several SOTA reinforcement learning algorithms in terms of sample efficiency on two robotic systems. Additionally, we perform ablation studies to evaluate the effectiveness of the proposed bootstrap and augmentation methods.
VI-A Training Setup
At the beginning of each training episode, a target state is randomly selected from a uniformly distributed range. For instance, in the robotic eye scenario, the target angles adhere to $\theta^{*}\sim\mathcal{U}[-15^{\circ},15^{\circ}]$ and $\psi^{*}\sim\mathcal{U}[-10^{\circ},10^{\circ}]$. The robotic agent strives to reach the target from the initial state throughout the episode. In this paper, simulation environments are all written in Python. Training and testing in simulation are conducted on a computer with an Intel Xeon Platinum 8383C CPU and an NVIDIA GeForce RTX 4090 GPU. The algorithms are implemented in PyTorch and trained using the Adam optimizer. The learning rate of DRL algorithms is set to $3\times 10^{-4}$ and the hidden layer size is fixed at $256$. The training hyperparameters $N$ and $M$ are empirically chosen based on the system’s complexity. Specifically, $N=2000$ and $M=250$ are used for the robotic eye, while $N=3500$ and $M=500$ are used for the robotic wrist. When the bootstrap method is not applied, random exploration is employed during the initial $N$ warm-up episodes. These hyperparameters are kept consistent across all comparison and ablation studies.
VI-B Comparison Results Against SOTA Methods
Several SOTA sample-efficient RL algorithms, designed for state-based observations and continuous action spaces, are selected for comparison. REDQ[34] integrates an ensemble of Q-functions with in-target minimization across a random subset of these functions. This allows for training with a high Update-To-Data (UTD) ratio, significantly improving sample efficiency. DroQ[35] equips the Q-functions with dropout connections and layer normalization techniques, maintaining a high UTD ratio. RLPD[36] leverages offline data during online learning by loading demonstrations into the replay buffer and oversampling them during training. X-QL[37] directly models soft optimal values and introduces a novel optimization framework based on the Gumbel distribution.
We trained the proposed SAC-BAR on two robotic systems and compared its performance with these SOTA approaches. Performance was measured by the online per-step average reward achieved by the agents, with each algorithm trained across four independent runs. Fig. 6(a) presents the comparative results between SAC-BAR and the selected SOTA RL methods. In the robotic eye system, SAC-BAR achieves performance comparable to (and slightly better than) DroQ and RLPD, while significantly outperforming REDQ and X-QL in terms of sample efficiency. In the robotic wrist system, SAC-BAR not only significantly outperforms RLPD and DroQ but also demonstrates moderate improvements in convergence speed over REDQ and X-QL. Notably, REDQ’s performance relies on a high UTD ratio and an ensemble of Q-functions, resulting in high computational cost; for instance, REDQ requires 18.2 hours for 3,500 training episodes, whereas SAC-BAR completes the same training in only 4.1 hours under our experimental setup. Moreover, X-QL exhibited significant instability during the early stages of training in both robotic systems. Overall, these results demonstrate that SAC-BAR is better suited than the existing SOTA RL methods for the string-type artificial muscle robots targeted in this study.
VI-C Ablation Study in Simulation
To evaluate the effectiveness of the proposed bootstrap and augmentation techniques, an ablation study was conducted on two robotic systems. Three variants of our proposed framework were tested as follows: (1) w/o A: This variant excludes the state vector augmentation method. (2) w/o B: This variant removes the bootstrap via PID method. (3) w/o AB: This baseline variant excludes both mechanisms. Each algorithm was trained and evaluated in simulation under randomized dynamics across four independent runs.
Training results of our proposed SAC-BAR, with and without bootstrap and augmentation, are presented in Fig. 6(b). As illustrated in the figures, our proposed algorithm exhibits the highest level of data efficiency in both robotic systems. Compared to the SAC algorithm trained solely with dynamics randomization (referred to as “w/o AB”), the integration of bootstrap and augmentation enhances data efficiency and expedites the training process. In the robotic eye system, the baseline algorithm achieves an average reward of $-0.10$ after 2,000 training episodes, whereas our approach achieves comparable performance in only approximately 750 episodes, resulting in a 2.67-fold increase in data efficiency compared to the original level. Similarly, in the robotic wrist system, the baseline algorithm achieves an average reward of $-1.61$ after 3,500 episodes, while ours achieves comparable performance in only about 1,200 episodes, leading to a 2.92-fold increase in data efficiency. Notably, in the case of the more complex robotic wrist system, we find the demonstrations provided by the PID controller yield a more significant acceleration effect during the early stages of training. Meanwhile, the augmentation method demonstrates a favorable impact on enhancing the training speed in both systems and throughout the entire learning process of the algorithms.
VII Real-World Experiments
This section presents the real-world experimental results of the proposed SAC-BAR on two testing robotic platforms. We first demonstrate the results of an ablation study conducted on the robotic wrist platform, which validate the effectiveness of the proposed dynamics randomization in enhancing control performance in real-world scenarios. Next, experimental results on the robotic eye are provided to show the generalizability of SAC-BAR across similar string-type artificial muscle-driven robotic systems. Additionally, we investigate the impact of the randomization range on control performance.
VII-A Ablation Study on Robotic Wrist
Figure 7: Comparison of experimental trajectories between the baseline and SAC-BAR. Target angles are set at ($\theta^{*}=2.5^{\circ},\phi^{*}=-7.5^{\circ}$). (a), (b) The system output angles ($\theta,\phi$) and corresponding actions $a_{t}$ of the baseline SAC algorithm. (c), (d) Those of our proposed SAC-BAR algorithm.
To evaluate the impact of each component on real-world control performances, three different control policies are tested on the robotic wrist platform, including the vanilla SAC algorithm, SAC with only randomization of muscles (SAC-R), and our proposed SAC-BAR. The parallel robotic wrist starts from the initial state with $\theta=\phi=0^{\circ}$ and attempts to reach and maintain the target angle set $(\theta^{*},\phi^{*})$. The action vector $a$ is computed by the trained actors at a time step of $t_{a}=0.5$ s, and each episode spans a duration of $t_{\text{total}}=20$ s. In Fig. 7(a), it is evident that without randomization of muscle dynamics, the vanilla-SAC struggles to achieve the desired orientation. While the agent may eventually reach the desired orientation at certain target angles, maintaining that target angle becomes a challenge, resulting in a considerable steady-state error. Conversely, as observed in Fig. 7(c), the proposed SAC-BAR agent achieves the target state in less than $10$ s and consistently remains within its neighborhood thereafter. As illustrated by the input action trajectories in Fig. 7(d), this success is attributed to the agent learning to maximize the action value during the rising phase and promptly reduce it upon approaching the target angles. Conversely, in Fig. 7(b), there exist serious oscillations in the action trajectories of vanilla-SAC during the initial $10$ s, leading to a prolonged rise time and sometimes even failure to reach the target.
Figure 8: Experimental results of the steady-state error tests in the parallel robotic wrist under selected control algorithms. (a), (b), (c) Heat maps illustrating the experimental results for the SAC, SAC-R, and SAC-BAR, respectively. The combination of target angles $(\theta^{*},\phi^{*})$ is systematically spaced at $2.5^{\circ}$ intervals within the range $(|\theta^{*}|,|\phi^{*}|)\leq 10^{\circ}$. The color bars in the visualization indicate the Euclidean distance norm of steady-state errors. (d) The box charts illustrating the medians, maxima, minima, and quartiles of the steady-state errors for the selected comparative control algorithms.
In order to comprehensively evaluate the control performance of the algorithms across different target angles, we designed a steady-state error field test for the robotic wrist. We selected a series of target angles $(\theta^{*},\phi^{*})$ spaced at $2.5^{\circ}$ intervals within the defined range $\theta^{*},\phi^{*}\in[-10^{\circ},10^{\circ}]$. Each agent was set to run for $25$ s, sequentially targeting each angle pair. The steady-state error is measured and averaged over the last $5$s window, determined by computing the Euclidean distance norm between the current and target angles, defined as $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\phi-\phi^{*})^{2}}$. The steady-state tracking errors for different targets of vanilla-SAC, SAC-R, and SAC-BAR are presented in Fig. 8(a), (b), and (c), respectively. The $\mathrm{x}$ and $\mathrm{y}$ axes therein represent $\theta^{*}$ and $\phi^{*}$, while the color bars represent the averaged steady-state errors. Following that, box charts are illustrated in Fig. 8(d). These figures demonstrate that vanilla-SAC agent underperforms across nearly the entire defined range, whereas both SAC-R and SAC-BAR show significantly improved control performance.
The steady-state errors averaged over the defined range, denoted as $\bar{e}_{\text{ss}}$, for three variants and the PID controller are presented in Table III. In real-world experiments, all three DRL-based controllers demonstrate superior performance compared to the traditional PID controller, reducing the tracking error by $11.6\%$, $35.9\%$, and $43.7\%$, respectively. Compared to the baseline SAC, SAC-R and SAC-BAR achieved reductions in the averaged tracking errors of $27.9\%$ and $36.3\%$, respectively. Specifically, our proposed SAC-BAR reaches the desired orientation at every target angle pair with an averaged error of $0.58^{\circ}$. Furthermore, as shown in Table III, the observed improvement in real-world performance can be attributed to overcoming the sim-to-real gap, as evidenced by the reduced disparity between simulated and actual results.
TABLE III: The experimental (Real) and simulation (Sim) results of steady state error using different control methods
$\bar{e}_{\text{ss}}$ (deg) Control Method
PID SAC SAC-R SAC-BAR
Real $1.03\pm 0.49$ $0.91\pm 0.46$ $0.66\pm 0.39$ $0.58\pm 0.37$
Sim $1.12\pm 0.47$ $0.46\pm 0.27$ $0.47\pm 0.33$ $0.40\pm 0.17$
The above experimental results indicate that the proposed randomization of muscle dynamics significantly enhances the real-world control performance in the robotic systems of interest, while bootstrap and augmentation techniques contribute relatively less to the improvement. We believe that the enhancement observed in SAC-BAR compared to SAC-R primarily stems from the improvement in training efficiency. With the incorporation of bootstrap and augmentation methods, SAC-BAR achieves more comprehensive training and higher averaged rewards in simulation after 3,500 training episodes. We believe this improvement is potentially more significant in scenarios where exploration is more challenging.
VII-B Learning Control of Robotic Eye
Figure 9: Experimental trajectories of the foveation test of the robotic eye. Target angles are set at $(\theta^{*}=-5^{\circ},\psi^{*}=7.5^{\circ})$ and $(\theta^{*}=10^{\circ},\psi^{*}=-5^{\circ})$, respectively. (a), (c) The system output angles $(\theta,\psi)$. (b), (d) The corresponding actions $a_{t}$ generated by the trained actor.
Figure 10: Experimental results on the steady-state tracking errors in robotic eye at eyery combination of $(\theta^{*},\psi^{*})$ target angles spaced at $2.5^{\circ}$ intervals for $(|\theta^{*}|,|\psi^{*}|)\leq 10^{\circ}$ where the color bar represents the Euclidean distance norm of steady-state error.
To demonstrate the applicability of the proposed learning control framework across diverse robotic systems driven by artificial muscle strings, this subsection presents experimental results on an additional platform, the 2-DOF robotic eye. We focus on the foveation control task in the experiment. The robotic eye, initialized at $\theta=\psi=0^{\circ}$ at ambient room temperature, is configured to foveate at various target angles $(\theta^{*},\psi^{*})$. Considering the system’s rise time, each testing episode lasts for $t_{\text{total}}=15$ s, with an action time step of $t_{a}=0.5$ s. The averaged Euclidean distance between the current and target angles within the last $5s$ window serves as the steady-state error metric, i.e., $e_{\text{ss}}=\sqrt{(\theta-\theta^{*})^{2}+(\psi-\psi^{*})^{2}}$.
Experimental trajectories with targets set at $(\theta^{*},\psi^{*})=(-5^{\circ},7.5^{\circ})$ and $(10^{\circ},-5^{\circ})$ are illustrated in Fig 9. These trajectories demonstrate that the measured angles rise to the target in approximately $5$ s, with negligible steady-state error. A similar field test is conducted with a series of target angles spaced at $2.5^{\circ}$ intervals within the defined range $|\theta^{*}|,|\psi^{*}|\leq 10^{\circ}$ to evaluate the overall control performance of the algorithm. The robotic eye is tested to foveate at each target in sequence. The results, illustrated in Fig. 10, reveal solid control performance, with steady-state errors consistently below $1^{\circ}$ for most target angles and an average of $0.56^{\circ}$.
In our prior research[20], we employed a DRL-based controller using the Deep Deterministic Policy Gradient algorithm[43], which was trained in simulation for 50,000 episodes. Additionally, a reduced-order state observer was used to estimate the real-time temperature of SCP muscles, achieving satisfactory control performance. In this paper, our proposed algorithm achieves comparable control performance with only 2,000 training episodes, representing just 1/25 of the original training time. Moreover, this is accomplished without the need to estimate real-time muscle temperature information under the condition of ambient temperature remaining relatively stable.
VII-C Randomization Ranges on Control Performance
TABLE IV: Comparison of control errors between models trained under different randomization ranges
$\bar{e}_{\text{ss}}$ (deg) Randomization Range
Small Original Large
Real $0.63\pm 0.49$ $0.58\pm 0.37$ $0.76\pm 0.43$
Sim $0.38\pm 0.17$ $0.40\pm 0.17$ $0.74\pm 0.41$
In designing dynamics randomization, maintaining a balance between randomization level and model fidelity is crucial. Excessive randomization leads to an overly conservative controller, which may over-generalize and fail to leverage the more predictable aspects of system dynamics. Conversely, insufficient randomization causes the controller to overfit to a narrow set of conditions, thus limiting its performance when deployed in real-world systems. To investigate this trade-off, we conducted an experiment in which our control framework was trained with two additional randomization ranges—“small” and “large.” The small range was defined as half of the original randomization range, while the large range was doubled. We evaluated the performance of controllers trained under these varying randomization ranges in both simulated and real-world environments on the robotic wrist platform. The results, as summarized in Table IV, indicate that the control performance experiences a significant decline when randomization range is excessively large. Conversely, the smaller randomization range impairs the controller’s generalization capability from simulation to reality. Overall, a balanced range proves essential for ensuring robust performance.
VIII Conclusion
This paper presents a learning-based control framework SAC-BAR designed for the string-type artificial muscle-driven robotic systems. To enhance training data efficiency and expedite convergence speed, we proposed two methods, namely bootstrap via PID and state-based data augmentation. Additionally, we incorporated randomization of artificial muscle dynamic parameters during training to address the sim-to-real gap for enhancing real-world control performance. Extensive experiments and ablation studies on two systems—a robotic eye driven and a parallel robotic wrist—validated the effectiveness and generalizability of the proposed control framework. In simulation, the designed bootstrap and augmentation methods reduced the required training episodes by approximately 60% compared to the baseline algorithm. Real-world experiments demonstrated that randomization of muscle dynamics improved the control performance of DRL-based controllers in targeted robotic systems, reducing the average error by 36.3%. Furthermore, our framework showed satisfactory performance on both robotic platforms, highlighting its potential applicability to similar artificial muscle robots.
Despite the strong performance of the proposed SAC-BAR, several limitations exist. The configuration of the reward function has a significant impact on experimental outcomes, often requiring empirical tuning to ensure optimal controller performance. The dynamic models of the artificial muscles and robotic wrist are simplified and do not fully represent their true dynamics. In future work, we plan to explore automated methods for reward function tuning, potentially incorporating adaptive or self-tuning mechanisms that respond dynamically to performance metrics. To better account for real-world variability, we will extend the randomization of muscle dynamics by incorporating high-fidelity models that capture nonlinearity and time-dependent factors, using data-driven or hybrid modeling approaches.